OpenStudy (anonymous):

Why could it be misleading for an owner of a salon to boast to her friends that the MEAN salary at the salon is $55,000?

OpenStudy (jamesj):

Because suppose the owner makes $200,000/year but the other 9 employees only make $38,889 a year. Then the mean would be \[ \frac{1}{10} ( 200,000 + 9 \cdot 38,889) = \frac{1}{10} 550,000 = 555,000 \] In other words, the mean doesn't tell you everything about the distribution of salaries and the fact that 90% of the employees make much less than that amount.

OpenStudy (jamesj):

*correction ... = 55,000

OpenStudy (jamesj):

Make sense?

OpenStudy (anonymous):

Yes, Thank you